2. Complete the table below for the outlet stream and show your working: Compound Molar mass(kg/kmol) Mass flow rate (kg/h) Mole flow rate (kmol/h) 02 CO2 CO H20 Total
Answer #1

Hii

Given – the mass of ethanol burnt = 276 kg/h

Molar mass of ethanol = CH3CH2OH=(12+3+12+2+16+1)=46 g/mole

Mass flow rate of 512 kg/hr

Molar mass of oxygen =32 g/mole

Moles = Mass in gram /molar mass in gram/mole

=molar flow rate of ethanol = 276000/46=6000 moles /hr

Molar flow rate of oxygen = 512000/32=16000 moles/hr

Combustion reaction of ethanol

C2H5OH+ 3O2= 2CO2+3H2O    selectivity 0.4

C2H5OH+ 2O2= 2CO+3H2O Selectivity = 0.6

As we are assuming complete combustion we can say that the moles of ethanol reacting via first and second process are 0.4:0.6

Assuming 1 hr of operation

C2H5OH+ 3O2= 2CO2+3H2O   

Moles of C2H2OH reacted = 0.4*6000= 2400 moles

Moles of O2 reacted = 3*2400= 7200 moles

Moles of CO2 produced = 2*2400= 4800 moles

Moles of H2O produced =3*2400=7200 moles

C2H5OH+ 2O2= 2CO+3H2O

Moles of C2H2OH reacted = 0.6*6000= 3600 moles

Moles of O2 reacted = 2*3600= 7200 moles

Moles of CO produced = 2*3600= 7200 moles

Moles of H2O produced =3*3600=10800 moles

Total moles in the product

CO2=4800 moles

Molar flow rate = 4800*44=211.2 Kg/hr

CO= 7200 moles

Molar flow rate = 7200*28=201.66 kg/hr

H20=10800 moles +7200 moles = 18000 moles

Molar flow rate = 18000*18=324 kg/hr

Remaining O2= 16000-7200-7200=1600 moles

Molar flow rate = 1600*32=51.2 Kg/hr

Molar mass of CO2= 12+32= 44 g/mole

Molar mass of CO= 12+16= 28 g/mole

Molar mass of H20= 2+16= 18 g/mole

Also g/mole= kg/kmole ( 1 kg=1000 g,1 kmoles =1000 moles)

Input mass flow rate

Let us complete the table now

Compound

Molar mass (kg/mole)

Mass flow rate (Kg/hr)

Molar flow rate (kmol/hr)

C2H5OH

46

276 input output-0

6

O2

32

512 input - 51.2 output

16 input – 1.6 output

CO2

44

211.2

4.8

CO

28

201.66

7.2

H2O

18

324

18

Total

788

   We can check the calculation here

Using law of conservation of mass

Mass in should be equal to mass out.

Mass in = 276+512 = 788 kg/hr

Mass out = 51.2+211.2+201.66+324= 787.5 kg/hr

We have taken some approximation while calculations so this little difference is coming.

Thank You